# RiverScope: High-Resolution River Masking Dataset

### v1

@@ 4.2 Baseline Methods
To evaluate (1) the effect of training on high-resolution satellite images and (2) the advantage of using pretrained models on the tasks, we compare results on several existing baselines. For (1), we compare performance against the same models trained on Sentinel data which have lower resolution. For (2), we compare on two models: a recent river width estimation model trained on PlanetScope data by Valman et al. (2024), and a widely used water segmentation algorithm NDWI (McFeeters 1996).
Sentinel models are trained on sampled Sentinel data from 2020-2021 with WorldCover (Van De Kerchove et al. 2021) labels for water segmentation. The same segmentation models, backbones, and pretraining methods from § 4.1 are also used to evaluate the effect of using low-resolution data on the downstream tasks. Multispectral data are used as input for all Sentinel models. We include more details in § A.4.
PlanetScope CNN (Valman et al. 2024) is a recent work that similarly labeled rivers from PlanetScope images. They use a randomly initialized custom CNN that was trained to segment water from satellite images, and was designed to take four bands as input.
PlanetScope NDWI (McFeeters 1996) uses the Normalized Difference Water Index (NDWI) to detect water from satellite images, which is widely used in remote sensing. NDWI is a value between -1 and 1 that uses the difference in the green and NIR reflectance values to detect water (Eq. 2), since water reflects almost no NIR. To find the threshold $t$ for water (i.e., $\text{NDWI}>t$), Otsu thresholding (Otsu et al. 1975) is applied.
@@ 5 Experiments
Figure 4: RiverScope trained models more accurately segment river pixels compared to Sentinel trained models. Each subplot shows the average F1 score improvement of a given segmentation model across multiple runs. For each architecture and pretraining combination, hatched bars represent the performance of Sentinel trained models, while solid bars represent the performance of RiverScope trained models. We show raw numbers in Table A3A2 (Appendix).
@@ 5.1 Water Segmentation
Using pretrained models improve segmentation performance. While PlanetScope CNN and PlanetScope NDWI also use high-resolution multispectral images, their performance is still lower compared to any of the RiverScope models that utilize transfer learning (Figure 4). Due to the rule-based nature of NDWI, it is sensitive to man-made land features and generally overestimates water boundaries which leads to falsely identifying pixels as water (Xu 2006) (see Table A2, Figure A3 (Appendix)). PlanetScope CNN tends to miss a significant number of water pixels, resulting in more false negatives and lower recall (see Figure A3, Figure A4 (Appendix)). Pretraining on large datasets such as ImageNet and SatlasPretrain can speed up the process of learning image features and result in better performance.
@@ 5.2 River Width Estimation
NDWI (McFeeters 1996) 194.6 51.0 160.2 176.3
CNN (Valman et al. 2024) 87.2 30.0 39.3 14.0
@@ 5.3 Future Directions
Sentinel predictions could potentially be further improved by removing outliers. While Sentinel has higher resolution than Landsat (10 m/px vs 30 m/px), the mean absolute error of Landsat (via SWORD) estimated widths is lower (Table 1). This may be attributed to the additional post-processing steps—such as basin-specific distribution fitting—applied to Landsat products, which can suppress outliers in the width estimates (Allen and Pavelsky 2018). Nonetheless, looking at median absolute error, a metric less sensitive to outliers, Sentinel widths still result in lower error compared to Landsat.
Combining SWOT and Sentinel predictions might offer a more robust approach for width estimation. Table 1 shows that SWOT and Sentinel achieve comparable median absolute error, but Sentinel underperforms in mean absolute error and bias. This bias stems from the high false positive rate in Sentinel’s water masks as detailed in § 5.1. Tying Sentinel widths to co-located SWOT measurements might help lessen this overestimation and reduce errors whenever optical observations are degraded by shadows, haze, clouds, or sunglint (Zhu and Helmer 2018). Incorporating SWOT’s KaRIn point cloud interferometry data—which provides geophysical data and radar echo power of samples—with Sentinel’s multispectral data may provide a physically grounded constraint on rivers that can reject optical/spectral outliers (Lin et al. 2020). Exploring this direction in future work may yield more reliable river width estimates when the optical data are limited or compromised.
@@ A.1 Limitations
Our dataset establishes a strong goundation with rivers of varying widths, primarily focused on the Northern Hemisphere due to the SWOT orbit. This provides an opportunity to expand the dataset to include more diversity of river systems in the Southern Hemisphere. The width labels in our work are also based on SWORD. While SWORD effectively defines single-channel reaches, future work could build on our work to address more complex morphologies, such as braided and multi-channel networks, which are challenging in river science.
@@
To achieve high precision, this study leveraged the high resolution of PlanetScope imagery. Its quality demonstrates the upper bound of what is possible with current technology. For future large-scale applications, our work can be used to evaluate the trade-off between purchasing more PlanetScope data and settling with less accuracy. Overcoming the coarser resolution of Sentinel and Landsat to achieve precise segmentation is a key challenge that our work helps to define. Furthermore, while all optical sensors are susceptible to atmospheric conditions like clouds and haze, the multimodal fusion approaches outlined below offer a robust path to mitigate these effects.
@@
Future research can directly address these points, building upon our contributions by (1) exploiting dense temporal stacks of Sentinel/Landsat to refine low-resolution segmentation, and (2) fusing SWOT’s precise height measurements with optical imagery. These approaches could effectively reduce false positives in water segmentation and provide a promising direction to explore multimodal learning for river science and water resource management.
A.2 Data Labeling Tool
@@
Figure A1 shows the data annotation tool for labeling PlanetScope images. We use the tool from Tangseng, Wu, and Yamaguchi (2017) as a starting point [^1] https://github.com/kyamagu/js-segment-annotator. The image on the upper left shows the RGB channels of the tile to be labeled, while the image on the upper right shows where annotators can label over the image. The SWORD nodes and reaches are overlaid on the upper left image as additional reference. The transparency and scale of the image on the upper right can be changed to see the finer details. The large image at the bottom shows the source PlanetScope tile, with the red box indicating the current tile being annotated. Additional information such as the reach identifier, latitude, longitude, and Google Maps link are also included in the upper left for the annotators’ reference.
@@
A.3 Additional RiverScope Details
@@
PlanetScope details. The PlanetScope data released as part of RiverScope were normalized using min-max normalization per image. That is, for each image, we get the maximum and the minimum values for normalization. Pixels with no data were filled with 0 values. We don’t use the minimum/maximum across the whole dataset since the raw data values are proprietary. The raw data can be purchased directly from PlanetLabs (PBC 2024) using the PlanetScope image IDs released in RiverScope.
@@
A.4 Sentinel Models
We use labeled data from WorldCover in 2020-2021 which has segmentation labels for land cover, including water segmentation labels. Only water labels are used for training Sentinel models. The training data was uniformly sampled on a global-scale, and is adopted from SatlasPretrain (Bastani et al. 2023). Figure A2 shows the distribution of the data with corresponding labels from WorldCover.
Sentinel trained models use 6 bands for training: red, blue, green, NIR, SWIR1, and SWIR2 (bands 2,3,4,8,11,12). This follows previous work (Daroya et al. 2025; Isikdogan, Bovik, and Passalacqua 2019) that predict masks on satellite images. Similar to RiverScope trained models, we use the Adam optimizer to train all models with the learning rate selected based on the performance on the validation set. The same binary cross entropy loss was also used. Segmentation performance was evaluated on the RiverScope test set, with the Sentinel predictions reprojected and upsampled to match the view and the resolution of the labels. Training time is about 12 hours on an NVIDIA L40S GPU for 50 epochs with a batch size of 32.
@@ A.5 Additional Results
PlanetScope NDWI tends to overestimate water boundaries and is sensitive to manmade structures. Table A2 shows the raw numbers for PlanetScope baselines CNN and NDWI. Despite the higher resolution satellite imagery, NDWI segments beyond the water boundaries, resulting in very precision. In addition, unlike other models that distinguish water from manmade structures (e.g., buildings), NDWI is sensitive to these classes. Table A4 shows that only NDWI has a non-zero false positive rate on ‘Built-up’, consistent with findings from Xu (2006).
@@
RN50 pretrained on ImageNet has competitive performance. Although the task requires reasoning on satellite images to discrimate between water and non-water pixels, our results show that pretraining on ImageNet1k (supervised) still results in superior performance. This is consistent with results from recent work that also show pretraining on ImageNet1k is still competitive compared to self-supervised remote sensing methods (Lahrichi et al. 2025).
@@
BCE loss results in best performance. Table A6 shows the performance of the different settings using different training losses. We explored several alternative loss functions commonly used in binary segmentation, including adaptive max-pooling (Isikdogan, Bovik, and Passalacqua 2019), Dice loss (Roy et al. 2017), and centerline-aware loss (Shit et al. 2021). However, in our experiments across multiple architectures and pre-training strategies, binary cross-entropy (BCE) consistently yielded the best average performance. While alternative losses can offer improvements in specific scenarios, BCE was the most robust and generalizable choice for our dataset and task.

### v2

@@ 4.2 Baseline Methods
To evaluate (1) the effect of training on high-resolution satellite images and (2) the advantage of using pretrained models on the tasks, we compare results on several existing baselines. For (1), we compare performance against the same models trained on Sentinel data which have lower resolution. For (2), we compare on two models: a recent river width estimation model trained on PlanetScope data by valman2024ai, and a widely used water segmentation algorithm NDWI (mcfeeters1996use).
Sentinel models are trained on sampled Sentinel data from 2020-2021 with WorldCover (van2021esa_worldcover) labels for water segmentation. The same segmentation models, backbones, and pretraining methods from § 4.1 are also used to evaluate the effect of using low-resolution data on the downstream tasks. Multispectral data are used as input for all Sentinel models. We include more details in the Appendix.
PlanetScope CNN (valman2024ai) is a recent work that similarly labeled rivers from PlanetScope images. They use a randomly initialized custom CNN that was trained to segment water from satellite images, and was designed to take four bands as input.
PlanetScope NDWI (mcfeeters1996use) uses the Normalized Difference Water Index (NDWI) to detect water from satellite images, which is widely used in remote sensing. NDWI is a value between -1 and 1 that uses the difference in the green and NIR reflectance values to detect water (Eq. 2), since water reflects almost no NIR. To find the threshold $t$ for water (i.e., $\text{NDWI}>t$), Otsu thresholding (otsu1975threshold) is applied.
@@ 5 Experiments
Figure 4: RiverScope trained models more accurately segment river pixels compared to Sentinel trained models. Each subplot shows the average F1 score improvement of a given segmentation model across multiple runs. For each architecture and pretraining combination, hatched bars represent the performance of Sentinel trained models, while solid bars represent the performance of RiverScope trained models. We show raw numbers in Table A4A3 (Appendix).
@@ 5.1 Water Segmentation
Using pretrained models improve segmentation performance. While PlanetScope CNN and PlanetScope NDWI also use high-resolution multispectral images, their performance is still lower compared to any of the RiverScope models that utilize transfer learning (Figure 4). Due to the rule-based nature of NDWI, it is sensitive to man-made land features and generally overestimates water boundaries which leads to falsely identifying pixels as water (xu2006modification) (see Table A3, Figure A4 (Appendix)). PlanetScope CNN tends to miss a significant number of water pixels, resulting in more false negatives and lower recall (see Figure A4, Figure A5 (Appendix)). Pretraining on large datasets such as ImageNet and SatlasPretrain can speed up the process of learning image features and result in better performance.
@@ 5.2 River Width Estimation
NDWI (mcfeeters1996use) 194.6 51.0 160.2 176.3
CNN (valman2024ai) 87.2 30.0 39.3 14.0
@@ 5.3 Future Directions
Sentinel predictions could potentially be further improved by removing outliers. While Sentinel has higher resolution than Landsat (10 m/px vs 30 m/px), the mean absolute error of Landsat (via SWORD) estimated widths is lower (Table 1). This may be attributed to the additional post-processing steps—such as basin-specific distribution fitting—applied to Landsat products, which can suppress outliers in the width estimates (allen2018global). Nonetheless, looking at median absolute error, a metric less sensitive to outliers, Sentinel widths still result in lower error compared to Landsat.
Combining SWOT and Sentinel predictions might offer a more robust approach for width estimation. Table 1 shows that SWOT and Sentinel achieve comparable median absolute error, but Sentinel underperforms in mean absolute error and bias. This bias stems from the high false positive rate in Sentinel’s water masks as detailed in § 5.1. Tying Sentinel widths to co-located SWOT measurements might help lessen this overestimation and reduce errors whenever optical observations are degraded by shadows, haze, clouds, or sunglint (zhu2018automatic). Incorporating SWOT’s KaRIn point cloud interferometry data—which provides geophysical data and radar echo power of samples—with Sentinel’s multispectral data may provide a physically grounded constraint on rivers that can reject optical/spectral outliers (lin2020global). Exploring this direction in future work may yield more reliable river width estimates when the optical data are limited or compromised.
6 Limitations
While our dataset presents rivers of varying widths, most rivers come from the Northern Hemisphere, providing an opportunity to expand to rivers in the Southern Hemisphere. Our width labels also rely on the river reaches defined by SWORD, which tend to simplify multiple or braided channels to a single channel, underestimating the widths in complex sections. The PlanetScope imagery used in our dataset has high resolution, but it is a commercial product—purchasing requirements limit access to additional data. Future large-scale applications can use our work to evaluate the trade-off between cost and accuracy. Additionally, all optical sensors remain vulnerable to haze, sunglint, and cloud cover. § A.1 (Appendix) discusses these limitations in detail.
@@ 7 Conclusion
Acknowledgements
We thank William Tooley for participating in the labeling of images. RD, TS, CG, and SM were supported in part by NASA grant 80NSSC22K1487. Additionally, RD and SM are supported by NSF grant 2329927, CG by NASA grant 80NSSC24K1646, and TR by a contract from the Jet Propulsion Laboratory.
@@ A.1 Limitations
Data distribution. Our dataset establishes a strong foundation with rivers of varying widths, primarily focused on the Northern Hemisphere due to the SWOT orbit. This provides an opportunity to expand the dataset to include more diversity of river systems in the Southern Hemisphere. The width labels in our work are also based on SWORD. While SWORD effectively defines single-channel reaches, future work could build on our work to address more complex morphologies, such as braided and multi-channel networks, which are challenging in river science.
@@
Cost for large-scale applications. To achieve high precision, this study leveraged the high resolution of PlanetScope imagery. Its quality demonstrates the upper bound of what is possible with current technology. For future large-scale applications, our work can be used to evaluate the trade-off between purchasing more PlanetScope data and settling with less accuracy. Overcoming the coarser resolution of Sentinel and Landsat to achieve precise segmentation is a key challenge that our work helps to define. Furthermore, while all optical sensors are susceptible to atmospheric conditions like clouds and haze, the multimodal fusion approaches outlined below offer a robust path to mitigate these effects.
Minimum detectable river width. The theoretical minimum detectable width is 3 meters (corresponding to a single pixel). However, based on our experience with previous computer vision models for river detection, we generally consider a river to be “detectable” when it spans at least three visible pixels across its width (i.e., approximately 9 meters wide for RiverScope).
Analyzing the river width estimation error as a function of the ground truth width, we find that the relative error generally increases with smaller rivers. Relative error is defined as $\frac{|y_{i}-\hat{y}_{i}|}{y_{i}}\times 100\%$ for a predicted width $\hat{y}_{i}$ and corresponding ground truth width $y_{i}$. Figure A1 summarizes this trend by showing the relative error across different ground truth river widths. The relative error is generally higher for smaller rivers, and lower for larger rivers. We found that while the model can detect narrow rivers, the relative error decreases significantly to 20% (9.8 meters MAE) for rivers around 45 meters wide (15 pixels).
A.2 Future Work
@@
Addressing limitations. Future research can directly address the limitations through building upon our contributions by (1) exploiting dense temporal stacks of Sentinel/Landsat to refine low-resolution segmentation, and (2) fusing SWOT’s precise height measurements with optical imagery. These approaches could effectively reduce false positives in water segmentation and provide a promising direction to explore multimodal learning for river science and water resource management.
Reservoir detection. Our work primarily focused on river segmentation, though some of our training data can include reservoirs since classify all water pixels. The model can often detect reservoir because of their visual similarity to rivers. However a more detailed evaluation would be needed to assess its accuracy in those cases.
Exploring other pretrained models. In this paper, we focused on widely used RGB and multispectral pretrained models. However, future work can look into models trained on high-resolution RGB-NIR similar to datasets fMoW (christie2018functional_fmow) and NAIP (usda_naip_imagery). Learning high-resolution features on these datasets could potentially further improve performance.
A.3 Data Labeling Tool
@@
Figure A2 shows the data annotation tool for labeling PlanetScope images. We use the tool from tangseng2017looking as a starting point [^1] https://github.com/kyamagu/js-segment-annotator. The image on the upper left shows the RGB channels of the tile to be labeled, while the image on the upper right shows where annotators can label over the image. The SWORD nodes and reaches are overlaid on the upper left image as additional reference. The transparency and scale of the image on the upper right can be changed to see the finer details. The large image at the bottom shows the source PlanetScope tile, with the red box indicating the current tile being annotated. Additional information such as the reach identifier, latitude, longitude, and Google Maps link are also included in the upper left for the annotators’ reference.
@@
Figure A1: Trend of relative error as the ground truth river width decreases. We generally see higher relative error for smaller rivers, and lower error for larger rivers.
@@
A.4 Additional RiverScope Details
@@
PlanetScope details. The PlanetScope data released as part of RiverScope were normalized using min-max normalization per image. That is, for each image, we get the maximum and the minimum values for normalization. Pixels with no data were filled with 0 values. We don’t use the minimum/maximum across the whole dataset since the raw data values are proprietary. The raw data can be purchased directly from PlanetLabs (planetlabs) using the PlanetScope image IDs released in RiverScope.
@@
Segmentation Model Backbone RGBR RGB+Random Linear
DeepLabv3 MV2 $93.6\pm 1.0$ $93.1\pm 1.1$ $93.7\pm 0.4$
RN50 $93.8\pm 0.2$ $93.5\pm 0.4$ $93.9\pm 0.4$
FPN MV2 $94.1\pm 0.1$ $93.7\pm 0.5$ $93.4\pm 0.7$
RN50 $93.1\pm 0.8$ $92.4\pm 1.3$ $93.5\pm 0.7$
UNet MV2 $92.8\pm 1.4$ $93.3\pm 0.4$ $94.0\pm 1.0$
RN50 $93.9\pm 0.3$ $92.8\pm 0.4$ $94.2\pm 0.4$
Average $93.6\pm 0.8$ $93.1\pm 0.8$ 93.8 $\pm$ 0.6
Table A2: Additional ways to handle NIR channel for pretrained RGB models. RGB+Random and RGBR use the original weights for the RGB channels in the first layer; they only differ in the way the NIR channel is handled. RGB+Random uses randomly initialized weights for NIR, while RGBR reuses the weights from the red channel for NIR. Linear still results in overall better performance.
@@
A.5 Sentinel Models
We use labeled data from WorldCover in 2020-2021 which has segmentation labels for land cover, including water segmentation labels. Only water labels are used for training Sentinel models. The training data was uniformly sampled on a global-scale, and is adopted from SatlasPretrain (bastani2023satlaspretrain). Figure A3 shows the distribution of the data with corresponding labels from WorldCover.
Sentinel trained models use 6 bands for training: red, blue, green, NIR, SWIR1, and SWIR2 (bands 2,3,4,8,11,12). This follows previous work (daroya2025improving; isikdogan2019seeing) that predict masks on satellite images. Similar to RiverScope trained models, we use the Adam optimizer to train all models with the learning rate selected based on the performance on the validation set. The same binary cross entropy loss was also used. Segmentation performance was evaluated on the RiverScope test set, with the Sentinel predictions reprojected and upsampled to match the view and the resolution of the labels. Training time is about 12 hours on an NVIDIA L40S GPU for 50 epochs with a batch size of 32.
@@ A.6 Additional Results
We also explore other ways to handle the NIR channel for pretrained RGB models such as RGB+Random and RGBR. RGB+Random keeps the original RGB pre-trained weights for the first layer of the model, and adds an additional channel with randomly initialized weights to handle the NIR channel. RGBR is similar to RGB+Random, but instead of randomly initializing the weights for the additional channel, the weights for the first channel are used. Scaling is applied to both methods to ensure the magnitude of the features do not change when fed to subsequent layers. Table A2 shows the results, where using a linear adaptor still resulted in overall better performance.
@@
PlanetScope NDWI tends to overestimate water boundaries and is sensitive to manmade structures. Table A3 shows the raw numbers for PlanetScope baselines CNN and NDWI. Despite the higher resolution satellite imagery, NDWI segments beyond the water boundaries, resulting in very precision. In addition, unlike other models that distinguish water from manmade structures (e.g., buildings), NDWI is sensitive to these classes. Table A5 shows that only NDWI has a non-zero false positive rate on ‘Built-up’, consistent with findings from xu2006modification.
@@
RN50 pretrained on ImageNet has competitive performance. Although the task requires reasoning on satellite images to discrimate between water and non-water pixels, our results show that pretraining on ImageNet1k (supervised) still results in superior performance. This is consistent with results from recent work that also show pretraining on ImageNet1k is still competitive compared to self-supervised remote sensing methods (lahrichi2025self).
@@
BCE loss results in best performance. Table A7 shows the performance of the different settings using different training losses. We explored several alternative loss functions commonly used in binary segmentation, including adaptive max-pooling (isikdogan2019seeing), Dice loss (roy2017relaynet), and centerline-aware loss (shit2021cldice). However, in our experiments across multiple architectures and pre-training strategies, binary cross-entropy (BCE) consistently yielded the best average performance. While alternative losses can offer improvements in specific scenarios, BCE was the most robust and generalizable choice for our dataset and task.